OpenStudy (anonymous):

If f(x)=√x+3 and g(x)=-7x+7 which statement is true? A.5is not the domain of (f o g) B. 5 is the domain of (f o g)

OpenStudy (anonymous):

an easy enough check, since \(g(5)=-7\times 5+7=-35+7=-28\)